OpenStudy (meshon):

The Doe family is ready to fill their new swimming pool. It can be filled in 12 hours if they use their own water hose, and in 30 hours if they use Mr. Jones', their neighbor's water hose. How long will the Doe's take to fill their pool if the neighbor's hose is used along with their own? * A. Assign a variable to represent the time required to fill the pool with both families' hoses and determine a rational expression that represent the part of the pool filled by both families' hoses in one hour. * B. Determine the rational expression that represents the time it takes the Doe's ho

OpenStudy (anonymous):

ANSWER: 8 4/7 hours (60/7) 1/12 + 1/30 = 1/x 5/60 + 2/60 = 1/x 7/60 = 1/x x = 60/7 <=>